# Projective reflection groups of finite covolume

### v1

@@ 1. Introduction
In this article, we characterize the reflection groups à la Vinberg that act with finite covolume on a properly convex domain, i.e. an open, convex and bounded subset $\Omega$ of an affine chart of a real projective space $\operatorname{\mathbb{P}}(V)$ of dimension $d\geq 1$. One can define a metric $d_{\Omega}$ on $\Omega$, called the Hilbert metric, that is invariant under the automorphisms of $\Omega$ (the linear transformations that leave $\Omega$ invariant). Our interest lies in the study of convex projective orbifolds, which are the quotients of properly convex domains by discrete subgroups of their automorphism groups.
@@
Many symmetric spaces admit convex projective structures; for example, the hyperbolic spaces, Euclidean spaces, and the symmetric space of $\operatorname{SL}(n,\mathbb{K})$ where $n\geq 3$ and $\mathbb{K}$ is the field of reals, complexes, or quaternions (however, the Hilbert metric differs from that of the symmetric space, except for the hyperbolic spaces). Moreover, $d_{\Omega}$ induces a Finsler metric and a measure on $\Omega$, and allows to ask whether a discrete subgroup $\Gamma$ of the automorphism group acts with finite covolume. In this case, we say that $\Gamma$ quasidivides $\Omega$, and that $\Omega$ is quasidivisible. In in fact $\Gamma$ acts cocompactly on $\Omega$, we say that $\Gamma$ divides $\Omega$ and that $\Omega$ is divisible. It is not difficult to see that cocompact lattices in $\operatorname{PO}(n,1)$ and $\operatorname{SL}(n,\mathbb{K})$ for $\mathbb{K}$ as above will provide many examples of divisible domains by considering their actions on the corresponding symmetric spaces (such examples are called symmetric).
@@
The first genuinely new examples of divisible domains in [KV67] used reflection groups in dimension $2$, and the theory was later greatly generalized in [Vin71]. A reflection group is a discrete subgroup $\Gamma$ of $\operatorname{SL^{\pm}}(V)$ that is generated by linear reflections along hyperplanes of $V$; it is abstractly isomorphic to a Coxeter group. Vinberg gave necessary and sufficient conditions on the reflections for $\Gamma$ to be a reflection group, and showed that an infinite reflection group always tiles a convex open subset of $\operatorname{\mathbb{P}}(V)$, called the Vinberg domain. When the Vinberg domain is properly convex, a reflection group can be seen as the group generated by chosen reflections along the facets of a properly convex polytope $P\subset\operatorname{\mathbb{P}}(V)$. The data of a properly convex polytope and reflections along its facets satisfying Vinberg’s conditions is called a Coxeter polytope. The Vinberg domain $\Omega_{P}$ of a Coxeter polytope $P$ does not have to be properly convex, but it is in most cases. When it is the case, we say that $P$ is of negative type.
@@
In the same article, Vinberg proved that the reflection group $\Gamma_{P}$ of a Coxeter polytope of negative type $P$ divides its Vinberg domain $\Omega_{P}$ if and only if it is perfect, i.e. for every vertex $v$ of $P$, the subgroup $\Gamma_{v}$ of $\Gamma_{P}$ generated by the reflections along the facets of $P$ containing $v$ is finite. He also showed that when $\Omega_{P}$ is symmetric, $\Gamma_{P}$ quasidivides $\Omega_{P}$ if and only if $P$ is quasiperfect, i.e. for all vertices $v$ of $P$, the group $\Gamma_{v}$ is either finite or conjugated to a maximal rank parabolic subgroup of $\operatorname{PO}(d,1)$.
@@
A similar condition was later introduced in [Mar17], as a regularity hypothesis: a Coxeter polytope $P$ is $2$-perfect if, for all vertices $v$ of $P$, the subgroup $\Gamma_{v}$ acts as a perfect reflection group on the space $V/\operatorname{Span}v$ (note that $v\in\operatorname{\mathbb{P}}(V)$ defines a $\Gamma_{v}$-invariant line $\operatorname{Span}v$ in $V$). Observe that perfect Coxeter polytopes are quasiperfect, and that quasiperfect Coxeter polytopes are $2$-perfect. Marquis showed in [Mar17] that if $P$ is a $2$-perfect Coxeter polytope of negative type, then $\Gamma_{P}$ quasidivides $\Omega_{P}$ if and only if $P$ is quasiperfect. Following a recent endeavor to investigate Vinberg theory without such regularity hypotheses (see [DGKLM25, ADLM25]), we generalize this result without any assumptions.
@@
Let $P$ be a Coxeter polytope of negative type. Then $P$ has finite volume in $\Omega_{P}$ if and only if $P$ is quasiperfect.
Prior to our work, the theoretic study of divisible domains had made significant progress in the sixties and the early two thousands (see [Ben60, Kos68, Vey70, Ben04, Ben03, Ben05, Ben06]). It appears that there are two types of divisible domains: either $\Omega$ is strictly convex, meaning that its boundary contains no non-trivial segments, in which case the group $\Gamma$ dividing is Gromov hyperbolic, or $\Omega$ is not strictly convex. In the strictly convex case, there exist examples that are non-symmetric divisible deformations of symmetric divisible domains [Gol90, JM87, Ben00] for $d\geq 2$ and that are not [Ben06a, Kap07] for $d\geq 4$. In the non-strictly convex case, there exist examples for $d\geq 3$ as well [Ben06, BDL18, CLM20, CLM22, LMR22, BV25].
@@
When $P$ is a quasiperfect Coxeter polytope of negative type, $\Omega_{P}/\Gamma_{P}$ has hyperbolic ends. Hence Theorem 1.1 shows that one cannot find examples of quasidivisible domains whose quotient presents generalized cusps (that are not hyperbolic cusps) using linear reflection groups.
Recent works in convex projective geometry have shown that in many geometrically interesting cases, there are infinitely many different invariant properly convex domains (see for instance [CLT18, DGK24, FIZ26]). However, it was known by [Vey70, Ben00] that in most cases, if a group $\Gamma$ divides a properly convex domain $\Omega\subset\operatorname{\mathbb{P}}(V)$, then $\Omega$ is the unique $\Gamma$-invariant properly convex domain in $\operatorname{\mathbb{P}}(V)$. The question remains open in the quasidivisible case, though Marquis had shown in [Mar17] that when $P\subset\operatorname{\mathbb{P}}(V)$ is a $2$-perfect Coxeter polytope of negative type, then $\Gamma_{P}$ preserves a unique properly convex domain in $\operatorname{\mathbb{P}}(V)$ (in most cases). Our proof of Theorem 1.1 also allows us to extend this theorem of Marquis by dropping the “$2$-perfectness” assumption.
Theorem 1.3 (see Theorem 7.1).
@@
More generally, we provide a characterization of the Coxeter polytopes of negative type for which the convex hull of the proximal limit set coincides with the Vinberg domain (see Corollary 7.3). This discussion is linked to the question of whether or not the proximal limit sets of quasidivisible domains fills their boundaries, as we explain at the end of Section 7. Indeed, Theorem 1.3 implies that if a Coxeter polytope $P$ of negative type is such that the proximal limit set of $\Gamma_{P}$ equals the boundary of its Vinberg domain $\Omega_{P}$, then $P$ is quasiperfect and $\Gamma_{P}$ is not virtually abelian (see Corollary 7.4). We conjecture that the reciprocal is true (see Conjecture 7.5).
@@
In Section 2, we give reminders about convex projective geometry, and in Sections 3 and 4, we list the definitions and tools we will need from Vinberg theory. In Section 5, we prove some fundamental lemmas. We then give in Sections 6 and 7 the proofs of Theorem 1.1 and 1.3 respectively.
Acknowledgements
@@ 2.1. Convex subsets of $\operatorname{\mathbb{P}}(V)$
Recall that in an affine space $\mathbb{A}$, the convex hull of a subset $X$ is the set of all linear combinations of points of $X$, and it is the smallest convex subset of $\mathbb{A}$ containing $X$. In projective space, there is a priori no good notion of convex hull since there are two projective segments between any two points. However, notice that if $\Omega\subset\operatorname{\mathbb{P}}(V)$ is a properly convex domain, then for any pair of distinct points $x,y\in\overline{\Omega}$, there is a unique projective segment contained in $\overline{\Omega}$ joining $x$ and $y$, corresponding to the segment between $x$ and $y$ in any affine chart of $\operatorname{\mathbb{P}}(V)$ containing $\Omega$. It follows that for any subset $X$ of $\overline{\Omega}$, the convex hull of $X$ in an affine chart $\mathbb{A}$ containing $\Omega$ does not depend on the choice of $\mathbb{A}$, so we may define $\operatorname{Conv}(X)$ to be the convex hull of $X$ in any affine chart containing $\Omega$.
@@ 2.2. Faces of a convex set
Let $C$ be a convex subset of $\operatorname{\mathbb{P}}(V)$. Recall that the span of $C$ is the smallest subspace of $V$ such that $C\subset\operatorname{\mathbb{P}}(\operatorname{Span}C)$. In fact, $C$ has non-empty relative interior in $\operatorname{\mathbb{P}}(\operatorname{Span}C)$. The dimension of $C$ is the dimension of $\operatorname{\mathbb{P}}(\operatorname{Span}C)$.
@@
Observe that a closed convex subset of $\operatorname{\mathbb{P}}(V)$ is the union of its faces. A priori, it may have uncountably many faces (e.g. an ellipsoid).
@@ 2.3. The Hilbert metric
where $x,y\in\Omega$ and $x^{\prime}$ and $y^{\prime}$ are the two intersections of $\partial\Omega$ with a projective line through $x$ and $y$, in such a way that $x^{\prime},x,y,y^{\prime}$ are aligned in that order (see Figure 1).
@@
Since the definition of $d_{\Omega}$ depends only on cross-ratios and $\Omega$ itself, the elements of $\operatorname{PGL}(V)$ that preserve $\Omega$ are isometries of $(\Omega,d_{\Omega})$, called automorphisms. For convenience, we prefer to work with subgroups of $\operatorname{SL^{\pm}}(V)\coloneqq\{g\in\operatorname{GL}(V)\mid\det g=\pm 1\}$, so we define the automorphism group by
@@ 3.1. Coxeter groups
A Coxeter system is a pair $(S,M)$ of a nonempty finite set $S$ and a symmetric matrix $M=(m_{st})_{s,t\in S}$ on $S$ with entries in $\mathbb{N}\cup\{\infty\}$ such that $m_{st}=1$ if and only if $s=t$ for all $s,t\in S$.
@@
The standard subgroups of $W=W_{S,M}$ are the subgroups $W_{T}$ of $W$ generated by a subset $T$ of $S$.
@@
Observe that $W$ is the direct product of the standard subgroups associated to the irreducible components of $S$. We say that $W$ is irreducible if it does not decompose as a direct product of non-trivial standard subgroups, or in other words, if $S$ has a unique irreducible component.
@@ 3.2. Vinberg’s theorem
Let $S$ be a finite set. We pick for each $s\in S$ a projective reflection $\sigma_{s}=\mathrm{Id}-\alpha_{s}\otimes v_{s}\in\operatorname{SL^{\pm}}(V)$ for some $(\alpha_{s},v_{s})\in V^{\ast}\times V$ such that $\alpha_{s}(v_{s})=2$, and we let
$\Delta\coloneqq\bigcap_{s\in S}\{\alpha_{s}\leq 0\}$
Assume furthermore that we chose the $\sigma_{s}$ so that $\Delta$ has non-empty interior.
Using Definition 2.2 in $V$ instead of $\operatorname{\mathbb{P}}(V)$, we can define the faces of $\Delta$. For any face $f$ of $\Delta$, we let $S_{f}\coloneqq\{s\in S\mid f\subset\ker\alpha_{s}\}$, and $\Gamma_{f}$ the subgroup of $\Gamma$ generated by the reflections $(\sigma_{s})_{s\in S_{f}}$.
@@
If it is the case, we say that $\Delta$ is the fundamental cone of $\Gamma$.
@@ A Cartan matrix on a finite set $S$ is a matrix $A=(A_{st})_{s,t\in S}$ satisfying:
for all $s,t\in S$. Then $(S,M)$ is a Coxeter system, so the Cartan matrix $A$ defines a Coxeter group $W_{A}\coloneqq W_{S,M}$ with generators in $S$. Notice that for any Coxeter group $W$, the Gram matrix $G_{W}$ of $W$ is a Cartan matrix such that $W_{G_{W}}=W$.
Theorem 3.8 ([Vin71, Thm. 1 & 2]).
Let $S$ be a finite set, and $(\sigma_{s})_{s\in S}$ be projective reflections in $\operatorname{SL^{\pm}}(V)$ such that $\Delta$ has non-empty interior. Then $\Gamma$ is a reflection group if and only if $A\coloneqq(\alpha_{s}(v_{t}))_{s,t\in S}$ is a Cartan matrix. In this case, we have:
@@
    $s\mapsto\sigma_{s}$ defines a discrete and faithful representation $\rho:W_{A}\to\operatorname{GL}(V)$ with image $\Gamma$;
@@
    $C\coloneqq\operatorname{int}\bigcup_{\gamma\in\Gamma}\gamma\Delta$ is a non-empty open convex cone of $V$;
@@
    $\Gamma$ acts properly discontinuously on $C$ with fundamental domain $\Delta\cap C$;
@@
    A face $f$ of $\Delta$ lies in $C$ if and only if $\Gamma_{f}$ is finite, otherwise $f\subset\partial C$;
If $\Gamma$ is a reflection group, the representation $\rho:W_{A}\to\operatorname{SL^{\pm}}(V)$ given by Vinberg theorem is called a Vinberg representation, and we say that $A$ is the Cartan matrix of $\Gamma$.

### v2

@@ 1. Introduction
In this article, we characterize the reflection groups à la Vinberg that act with finite covolume on a properly convex domain, i.e. an open, convex and bounded subset $\Omega$ of an affine chart of a real projective space $\operatorname{\mathbb{P}}(V)$ of dimension $d\geq 1$. One can define a metric $d_{\Omega}$ on $\Omega$, called the Hilbert metric, that is invariant under the automorphisms of $\Omega$, i.e. the linear transformations that leave $\Omega$ invariant. Our interest lies in the study of convex projective orbifolds, which are the quotients of properly convex domains by discrete subgroups of their automorphism groups. Many symmetric spaces admit convex projective structures, such as the real hyperbolic spaces, Euclidean spaces, and the symmetric spaces of $\operatorname{SL}(n,\mathbb{K})$ for $n\geq 3$ and $\mathbb{K}$ the field of the real, complex, or quaternionic numbers (note however that the Hilbert metric differs from that of the symmetric space, except for the real hyperbolic spaces).
@@
The Hilbert metric $d_{\Omega}$ of a properly convex domain $\Omega$ induces a Finsler metric and a Borel measure on $\Omega$, both invariant under the automorphisms of $\Omega$. This allows one to ask whether a discrete subgroup $\Gamma$ of the automorphism group acts with finite covolume. If this is the case, we say that $\Gamma$ quasidivides $\Omega$, and that $\Omega$ is quasidivisible. If in fact $\Gamma$ acts cocompactly on $\Omega$, we say that $\Gamma$ divides $\Omega$ and that $\Omega$ is divisible. It is not difficult to see that (cocompact) lattices of the Lie groups whose symmetric space admit a convex projective structure easily provide examples of (quasi)divisible domains (such examples are called symmetric).
@@
The first genuinely new examples of divisible domains in [KV67] used reflection groups in dimension $2$, and the theory was later greatly generalized in [Vin71]. A reflection group is a discrete subgroup $\Gamma$ of $\operatorname{SL^{\pm}}(V)$ that is generated by linear reflections along hyperplanes of $V$; it is abstractly isomorphic to a Coxeter group. Vinberg gave in [Vin71] necessary and sufficient conditions on the reflections for $\Gamma$ to be a reflection group, and showed that an infinite reflection group always tiles a convex open subset of $\operatorname{\mathbb{P}}(V)$, called the Vinberg domain. When the Vinberg domain is properly convex, a reflection group can be seen as the group generated by chosen reflections along the facets of a properly convex polytope $P\subset\operatorname{\mathbb{P}}(V)$. The data of a properly convex polytope and reflections along its facets satisfying Vinberg’s conditions is called a Coxeter polytope. The Vinberg domain $\Omega_{P}$ of a Coxeter polytope $P$ does not have to be properly convex, but it is in most cases. When it is the case, we say that $P$ is of negative type.
@@
In the same article, Vinberg proved that the reflection group $\Gamma_{P}$ of a Coxeter polytope of negative type $P$ divides its Vinberg domain $\Omega_{P}$ if and only if $P$ is perfect, i.e. for every vertex $v$ of $P$, the subgroup $\Gamma_{v}$ of $\Gamma_{P}$ generated by the reflections along the facets of $P$ containing $v$ is finite. He also showed that when $\Omega_{P}$ is symmetric, $\Gamma_{P}$ quasidivides $\Omega_{P}$ if and only if $P$ is quasiperfect, i.e. for all vertices $v$ of $P$, the group $\Gamma_{v}$ is either finite or conjugated to a maximal rank parabolic subgroup of $\operatorname{PO}(d,1)$.
@@
A similar condition was later introduced in [Mar17], as a regularity hypothesis: a Coxeter polytope $P$ is $2$-perfect if, for all vertices $v$ of $P$, the subgroup $\Gamma_{v}$ acts as a perfect reflection group on the space $V/\operatorname{Span}v$ (note that $v\in\operatorname{\mathbb{P}}(V)$ defines a $\Gamma_{v}$-invariant line $\operatorname{Span}v$ in $V$). Observe that perfect Coxeter polytopes are quasiperfect, and that quasiperfect Coxeter polytopes are $2$-perfect. Marquis showed in [Mar17] that if $P$ is a $2$-perfect Coxeter polytope of negative type, then $\Gamma_{P}$ quasidivides $\Omega_{P}$ if and only if $P$ is quasiperfect. Following a recent endeavor to investigate Vinberg theory without such regularity hypotheses (see [DGKLM25, ADLM25]), we generalize this result with minimal assumptions.
@@
Let $P$ be a Coxeter polytope of negative type. Then $\Gamma_{P}$ quasidivides $\Omega_{P}$ if and only if $P$ is quasiperfect.
Prior to our work, the theoretic study of divisible domains had made significant progress in the sixties and the early two thousands (see [Ben60, Kos68, Vey70, Ben04, Ben03, Ben05, Ben06]). It appears that there are two types of divisible domains exhibiting different behaviors: either $\Omega$ is strictly convex, meaning that its boundary contains no non-trivial segments, in which case the group $\Gamma$ dividing $\Omega$ is Gromov hyperbolic, or $\Omega$ is not strictly convex. In the strictly convex case, there exist examples that are non-symmetric divisible deformations of symmetric divisible domains [Gol90, JM87, Ben00] for $d\geq 2$ and that are not [Ben06a, Kap07] for $d\geq 4$. In the non-strictly convex case, there are examples for $d\geq 3$ as well [Ben06, BDL18, CLM20, CLM22, LMR22, BV25].
@@
When $P$ is a quasiperfect Coxeter polytope of negative type, $\Omega_{P}/\Gamma_{P}$ has hyperbolic ends. Hence, Theorem 1.1 shows that one cannot find examples of quasidivisible domains whose quotient presents generalized cusps (that are not hyperbolic cusps) using linear reflection groups.
Recent works in convex projective geometry have shown that in many geometrically interesting cases, there are infinitely many different invariant properly convex domains (see for instance [CLT18, DGK24, FIZ26]). However, it was known by [Vey70, Ben00] that in most cases, if a group $\Gamma$ divides a properly convex domain $\Omega\subset\operatorname{\mathbb{P}}(V)$, then $\Omega$ is the unique $\Gamma$-invariant properly convex domain in $\operatorname{\mathbb{P}}(V)$. The question remains open in the quasidivisible case, though Marquis had shown in [Mar17] that when $P\subset\operatorname{\mathbb{P}}(V)$ is a $2$-perfect Coxeter polytope of negative type, $\Gamma_{P}$ preserves a unique properly convex domain in $\operatorname{\mathbb{P}}(V)$ if and only if $P$ is quasiperfect. Theorem 1.1 allows us to extend this theorem of Marquis by dropping the “$2$-perfectness” assumption.
Theorem 1.3.
@@
More generally, we provide a characterization of the Coxeter polytopes of negative type for which the convex hull of the proximal limit set coincides with the Vinberg domain (see Theorem 7.1). This discussion is linked to the question of whether or not the proximal limit set of quasidivisible domains fills their boundaries, as we explain at the end of Section 7. Indeed, Theorem 1.3 implies that if a Coxeter polytope $P$ of negative type is such that the proximal limit set of $\Gamma_{P}$ equals the boundary of its Vinberg domain $\Omega_{P}$, then $P$ is quasiperfect and $\Gamma_{P}$ is not virtually abelian (see Corollary 7.5). We conjecture that the reciprocal is true (see Conjecture 7.6).
@@
In Section 2, we give reminders about convex projective geometry, and in Sections 3 and 4, we list the definitions and tools we will need from Vinberg theory. In Section 5, we prove elementary results necessary for the proof of our main theorems. We then give in Sections 6 and 7 the proofs of Theorems 1.1 and 1.3 respectively.
Acknowledgments
@@ 2.1. Convex subsets of $\operatorname{\mathbb{P}}(V)$
Recall that in an affine space $\mathbb{A}$, the convex hull of a subset $X$ is the set of all convex combinations of points of $X$, and it is the smallest convex subset of $\mathbb{A}$ containing $X$. In projective space, there is a priori no good notion of convex hull since there are two projective segments between any two points. However, notice that if $\Omega\subset\operatorname{\mathbb{P}}(V)$ is a properly convex domain, then for any pair of distinct points $x,y\in\overline{\Omega}$, there is a unique projective segment contained in $\overline{\Omega}$ joining $x$ and $y$, corresponding to the segment between $x$ and $y$ in any affine chart of $\operatorname{\mathbb{P}}(V)$ containing $\Omega$. It follows that for any subset $X$ of $\overline{\Omega}$, the convex hull of $X$ in an affine chart $\mathbb{A}$ containing $\Omega$ does not depend on the choice of $\mathbb{A}$, so we may define $\operatorname{Conv}(X)$ to be the convex hull of $X$ in any affine chart containing $\Omega$.
@@ 2.2. Faces of a convex set
Let $C$ be a convex subset of $\operatorname{\mathbb{P}}(V)$. Recall that the span of $C$, denoted by $\operatorname{Span}C$, is the smallest subspace of $V$ such that $C\subset\operatorname{\mathbb{P}}(\operatorname{Span}C)$. In fact, $C$ has non-empty relative interior in $\operatorname{\mathbb{P}}(\operatorname{Span}C)$. The dimension of $C$ is the dimension of $\operatorname{\mathbb{P}}(\operatorname{Span}C)$.
@@
Observe that a closed convex subset of $\operatorname{\mathbb{P}}(V)$ is the union of its faces. A priori, it may have uncountably many faces (e.g. an ellipsoid). We order the faces of a closed convex subset by the inclusion of the closures: $f$ is smaller than $g$ if $\overline{f}\subset\overline{g}$.
@@ 2.3. The Hilbert metric
for $x\neq y\in\Omega$, where $x^{\prime}$ and $y^{\prime}$ are the intersections of $\partial\Omega$ with the projective line through $x$ and $y$, in such a way that $x^{\prime},x,y,y^{\prime}$ are aligned in that order (see Figure 1).
@@
Since the definition of $d_{\Omega}$ depends only on cross-ratios and $\Omega$ itself, the elements of $\operatorname{PGL}(V)$ that preserve $\Omega$ are isometries of $(\Omega,d_{\Omega})$, called automorphisms. For convenience, we prefer to work with subgroups of
$\operatorname{SL^{\pm}}(V)\coloneqq\{g\in\operatorname{GL}(V)\mid\det g=\pm 1\},$
so we define the automorphism group of $\Omega$ by
@@ 3.1. Coxeter groups
A Coxeter system is a pair $(S,M)$ of a non-empty finite set $S$ and a symmetric matrix $M=(m_{st})_{s,t\in S}$ on $S$ with entries in $\mathbb{N}\cup\{\infty\}$ such that $m_{st}=1$ if and only if $s=t$ for all $s,t\in S$.
@@
A standard subgroup of $W=W_{S,M}$ is the subgroup $W_{T}$ of $W$ generated by a subset $T$ of $S$.
@@
Observe that $W$ is the direct product of the standard subgroups associated to the irreducible components of $S$. We say that $W$ is irreducible if it does not decompose as a direct product of non-trivial standard subgroups, or in other words, if $S$ has a unique irreducible component. $W$ is said to be reducible if it is not irreducible.
@@ 3.2. Vinberg’s theorem
Let $S$ be a finite set. We pick a projective reflection $\sigma_{s}=\mathrm{Id}-\alpha_{s}\otimes v_{s}\in\operatorname{SL^{\pm}}(V)$ for each $s\in S$, where $(\alpha_{s},v_{s})\in V^{\ast}\times V$ satisfies $\alpha_{s}(v_{s})=2$, and we let
$\Delta\coloneqq\bigcap_{s\in S}\{v\in V\mid\alpha_{s}(v)\leq 0\}$
Assume furthermore that we chose $(\sigma_{s})_{s\in S}$ so that $\Delta$ has non-empty interior.
@@
If this is the case, we say that $\Delta$ is the fundamental cone of $\Gamma$.
Using Definition 2.2 in $V$ instead of $\operatorname{\mathbb{P}}(V)$, we can define the faces of $\Delta$. For any face $f$ of $\Delta$, we let $S_{f}\coloneqq\{s\in S\mid f\subset\ker\alpha_{s}\}$, and we denote by $\Gamma_{f}$ the subgroup of $\Gamma$ generated by the reflections $(\sigma_{s})_{s\in S_{f}}$.
@@ A Cartan matrix on a finite set $S$ is a matrix $A=(A_{st})_{s,t\in S}$ satisfying:
for all $s,t\in S$. Then $(S,M)$ is a Coxeter system, so the Cartan matrix $A$ defines a Coxeter group $W_{A}\coloneqq W_{S,M}$. Notice that for any Coxeter group $W$, the Gram matrix $G_{W}$ of $W$ is a Cartan matrix that satisfies $W_{G_{W}}=W$.
Fact 3.8 ([Vin71, Thm. 1 & 2]).
Let $S$ be a finite set, and $(\sigma_{s})_{s\in S}$ be projective reflections in $\operatorname{SL^{\pm}}(V)$ such that $\Delta$ has non-empty interior. Then $\Gamma$ is a reflection group if and only if $A\coloneqq(\alpha_{s}(v_{t}))_{s,t\in S}$ is a Cartan matrix. In this case:
@@
    $s\mapsto\sigma_{s}$ defines a discrete and faithful representation $\rho:W_{A}\to\operatorname{GL}(V)$ with image $\Gamma$.
@@
    $C\coloneqq\operatorname{int}\bigcup_{\gamma\in\Gamma}\gamma\Delta$ is a non-empty open convex cone of $V$.
@@
    $\Gamma$ acts properly discontinuously on $C$ with fundamental domain $\Delta\cap C$.
@@
    A face $f$ of $\Delta$ lies in $C$ if and only if $\Gamma_{f}$ is finite; otherwise $f\subset\partial C$.
If $\Gamma$ is a reflection group, then the representation $\rho:W_{A}\to\operatorname{SL^{\pm}}(V)$ given by Fact 3.8 is called a Vinberg representation, and we say that $A$ is the Cartan matrix of $\Gamma$.